[Brahms-l] Shift report 20050215 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Tue Feb 15 2005 - 00:03:47 EST
Shifter leader was hy
Shifters were 


1. Came with beam studies ongoing, got beam for physics at around 17:35,  dumped scheduled for 0:40am.

2. Setting 4A3450 + 40A1050
RunNo  -  Trig2    -  Trig3
13185  -    197k   -   119k
13188  -    ~70k   -   ~80k 
(for r13188, got some RunError related to "spool output stream missing")
13190  -    127k   -   215k
13191  -    191k   -   172k
13192  -    stopped because of IOServer problem
13193  -    ongoing
3. None found

Shifter: hy

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue Feb 15 00:04:13 2005

This archive was generated by hypermail 2.1.8 : Tue Feb 15 2005 - 00:04:27 EST